Find the value of following 1. 4cot45O – sec60O +sin30O
1
Expert's answer
2020-05-25T20:11:16-0400

1. 4cot45°- sec60° +sin30°


4cot45°=4×1=4;

sec 60°=1/cos 60°=1/0.5=2;

sin 30°= 1/2;

4cot45°- sec60° +sin30°= 4-2+1/2=2+1/2=5/2=2.5

Answer: 4cot45°- sec60° +sin30°= 5/2